# CounterCurate: Enhancing Physical and Semantic Visio-Linguistic Compositional Reasoning via Counterfactual Examples

### v2

@@ Abstract
We first spotlight the near-chance performance of multimodal models like CLIP and LLaVA in physically grounded compositional reasoning. We then apply simple data augmentation using grounded image generation model GLIGEN to generate fine-tuning data, resulting in significant performance improvements: +33% and +37% for CLIP and LLaVA, respectively, on our newly curated Flickr30k-Positions benchmark. Moreover, we exploit the capabilities of high-performing text generation and image generation models, specifically GPT-4V and DALLE-3, to curate challenging semantic counterfactuals, thereby further enhancing compositional reasoning capabilities on benchmarks such as SugarCrepe, where CounterCurate outperforms GPT-4V.
To facilitate future research, we will release our code, dataset, benchmark, and checkpoints at https://countercurate.github.io/.
@@
[…]
@@ 1 Introduction
At present, the majority of research focuses on semantic compositional reasoning, leaving another fundamental problem, physically grounded compositional reasoning, under-explored. This includes tasks such as counting and distinguishing left/right and up/down positions between objects. For example, in Figure 1, GPT-4V OpenAI (2023c) notably made the wrong choice even though it identified all objects in question, demonstrating capable semantical yet weak physical compositional reasoning.
@@
With such an observation, it is not surprising to find that both contrastive models like CLIP Radford et al. (2021) and generative models like LLaVA Liu et al. (2023b) deliver near random chance performance on our newly curated benchmarks. We hypothesize that modern LMMs are largely oblivious to positional differences. To address this, we generate counterfactual image-text pairs using both simple methods, such as horizontal flipping for left/right distinctions, and leveraging a grounded image generation model, GLIGEN Li et al. (2023), to accurately replace/remove objects of interest for up/down and counting tasks. Our approach shows significant improvements such as 33% for CLIP and 37% for LLaVA on our Flickr30k-Positions benchmark.
Existing methods also do not fully utilize the capabilities of powerful generative models when creating semantic counterfactual fine-tuning data. We argue that the key to enhancing compositional reasoning capabilities lies in the careful curation of accurate and sufficiently challenging image and text counterfactuals. For example, augmented negative captions with linguistic rules used by recent fine-tuning approaches (e.g., Yuksekgonul et al., 2023; Le et al., 2023) could unintentionally follow unnatural language distributions, which are easily distinguishable by a text-only model without any image evidence (Lin et al., 2023).
@@ 2.2 Compositional Reasoning
NegCLIP Yuksekgonul et al. (2023) attempts to improve compositional reasoning via fine-tuning CLIP with perturbed captions as the negative text samples in contrastive learning. However, such permuted captions can be easily detected by a text-only transformer Lin et al. (2023), demonstrating that a more effective approach is needed to further enhance compositional reasoning. DAC Doveh et al. (2024) leverages dense and aligned captions for better positive captions. SGVL Herzig et al. (2023) utilizes scene graph to fine-tune the pretrained vision-language model. TSVLC Doveh et al. (2023) leverages language models to generate positive and negative captions. COCO-Counterfactual Le et al. (2023) further utilizes both negative images and negative captions to fine-tune CLIP via a rule-based data curation pipeline.
Moreover, previous works mainly consider composition reasoning with semantically different concepts rather than physically grounded relationships such as object counting and distinguishing left/right and above/below. Though a recent work Paiss et al. (2023) improved CLIP’s counting capabilities, they neither addressed compositional reasoning nor leveraged negative images. In our paper, we systematically improve both semantic and physical compositional reasoning by leveraging the most capable generative models including GPT-4V OpenAI (2023c) and DALLE-3 Betker et al. (2023).
@@ 2.3 Counterfactual Reasoning
Figure 3: Fine-tuning different types of large multimodal models with CounterCurate. Our pipeline can enhance both contrastive learning models and generative models by augmenting vanilla image-caption pairs with curated negative images and captions. Specifically, our counterfactual image-caption pairs (a) provide auxiliary contrastive loss for models like CLIP, where positive contrastive units in the similarity matrix are painted as blue/red and negative ones are painted as white, and (b) can be naturally integrated into the original next-token prediction loss in text generation models such as LLaVA.
Counterfactual reasoning Morgan and Winship (2015) refers to the process of imagining “what if” scenarios by modifying the input data. In the context of visio-linguistics scenarios, this involves curating negative images and captions by manipulating the original data and observing how it affects the outcome. This helps models understand cause and effect and predict outcomes in situations they’ve never seen before. The key to counterfactual reasoning is curating meaningful and hard enough examples. COCO-Counterfactual Le et al. (2023) explores simple linguistic rules to generate negative captions and uses an image editing model Prompt2Prompt Hertz et al. (2022) to produce negative images Hertz et al. (2022). Prompt2Prompt is less desirable in understanding complex language instructions, making the generated counterfactuals less reliable and challenging.
In this paper, we collect the most challenging negative image and text examples by leveraging the most capable language models GPT-4 OpenAI (2023b), text-to-image generation model DALLE-3 Betker et al. (2023), and GLIGEN Li et al. (2023), significantly improving our model’s reasoning. Utilizing these hard negatives, our model can better grasp the nuances of language and vision, enhancing its performance on tasks that require a sophisticated understanding of the world.
@@ 3 Approach
In Sec. 3.1, we first justify our selecting Flickr30k Entities dataset for both benchmarking and fine-tuning. Further in Sec. 3.2 and 3.3, we demonstrate the near-chance performance of multimodal models like CLIP and LLaVA on physically grounded compositional reasoning tasks, including positional understanding and counting, and improve their performance by generating negative image-text pairs via data augmentation and grounded image inpainting. Finally, we introduce using capable text and image generation models to improve models’ semantic compositional reasoning in Sec. 3.4.
@@ 3.1 Necessity of Grounded Image Captions
Previous approaches Yuksekgonul et al. (2023); Hsieh et al. (2023) adopt global image captioned datasets such as COCO-Captions Chen et al. (2015) and craft negative captions to either benchmark or improve compositional reasoning. This method, however, lacks regional information important to physically grounded reasoning, such as determining the relative position of two objects. Therefore, we leverage Flickr30k Entities Plummer et al. (2016), a grounded image captioning dataset to both curate negative image-caption pairs and benchmark models.
@@ 3.2 Improving Positional Understanding
in Flickr30k Entities Plummer et al. (2016) for each object $j$ in Image $I$ that only has one bounding box, where box $B_{j}$’s upper-left and lower-right corners are at $(x_{j1},y_{j1})$ and $(x_{j2},y_{j2})$, respectively. For each pair of objects $\{a,b\}\subseteq I$, we identify the positional object pairs via the formula:
[…]
@@
We also prompt text-only GPT4 OpenAI (2023b) to rewrite the vanilla negative prompt $a$ is above/below $b$ to better guide image generation models. For example, the caption “a street is above young child" is very unnatural, while GPT-expanded caption “the street stretches out above a young child, elevated by a bridge” is much better. For detailed prompt engineering, see Appendix B.
@@
Furthermore, we include negative images in our dataset to curate Flickr30k-Positions. For the left-right pairs, we apply a simple data augmentation method: flip the original images horizontally. For the above-below pairs, however, flipping vertically without context would lead to unreasonable images. To address this issue, we utilize GLIGEN Li et al. (2023), an object-level text-to-image generation model that can ground text prompts with bounding boxes. In particular, for each above-below object pair $(a,b)\in I$, we swap the centers of the bounding boxes $B_{a},B_{b}$ while maintaining their widths and heights, ensuring that the generated images are natural while preserving the objects’ sizes and aspect-ratios as demonstrated in Figure 4. The new bounding boxes and prompts are fed to GLIGEN to generate the corresponding negative image $I^{\prime}$.
@@ 3.3 Improving Object Counting
If both objects do not spatially overlap with other objects, we generate the corresponding negative caption by decrementing the one with more objects, $Q$, and incrementing the one with fewer, $P$; for example, $C^{\prime}=$“there are four cats and three dogs”. This enforces a hard counterfactual format as the numbers in the positives are reversed in the negatives.
In most cases, however, objects in an image do overlap, and we cannot ascertain that the increment/decrement rule still generates hard counterfactuals. For example, if a dog overlaps with a cat, and we re-use the rule above to remove one dog and add one cat, we end up with “three cats and three dogs”. In this case, we can only remove either one of $P$ or one of $Q$, for example, the dog, and all objects it overlaps with. An example of the resulting negative caption is $C^{\prime}=$“there are two cats and three dogs.”
@@
In the case where objects do not overlap, we simply leverage GLIGEN Li et al. (2023) to inpaint a new object of the incremented category $P$ at the bounding box of one of the objects from the decremented category $Q$. Otherwise, we remove the target objects using GLIGEN to replace it with an object from a generic category. We simply use the plant category for replacement, as plants can appear in almost any natural setting, let it be outdoor or indoor, and we find that GLIGEN does well in generating it according to each image $I$’s background during inpainting for each corresponding negative image $I^{\prime}$. More details in Appendix C.
@@ 3.4 Improving Semantic Compositional Reasoning
Existing works Yuksekgonul et al. (2023); Le et al. (2023) use a rule-based approach to generate negative images/captions for semantic compositional reasoning, resulting in less desirable negatives such as uncommon negative captions including “A dog with blue fur" from “A dog with black fur". We introduce Flickr30k-Attributes as in Figure 2 (c), which utilizes GPT-4V OpenAI (2023c) to generate hard negative captions and leverages DALLE-3 Betker et al. (2023) for generating corresponding high-quality negative images. We extract the first and most detailed caption $C$ for each image $I$ from the original Flickr30k dataset Young et al. (2014).
@@
To generate high-quality negative captions, we leverage GPT-4V and feed both positive images $I$ and captions $C$ to generate negative captions $C^{\prime}$. To let GPT-4V know the objects’ locations, we overlay bounding boxes $B$ onto the original positive image Cai et al. (2023) and also feed this annotated image into GPT-4V. We prompt GPT-4V to generate three kinds of captions, including (i) changing a noun, (ii) changing an adjective, and (iii) swapping any of the two phrases, nouns, adjectives, etc. For example, given the original caption $C$ “a man wearing a black shirt and blue jeans", GPT-4V can generate corresponding negatives $C^{\prime}$: (i) “a man wearing a black jacket and blue jeans", (ii) “a man wearing a black shirt and red jeans", and (iii) “a man wearing a blue shirt and black jeans". Detailed prompts to GPT-4V are shown in Appendix A.
@@ 3.5 Fine-tuning Methodology
As demonstrated in Figure 3, CounterCurate can be used to fine-tune both contrastive Radford et al. (2021) and generative Liu et al. (2023b) multimodal models, which significantly improve their performance. To fine-tune a contrastive model like CLIP, we sample $N$ positive image-text pairs and the corresponding negative image-text pairs to form a batch $(C,I,C^{\prime},I^{\prime})$ as shown in Figure 3 (a). We refer to this method as Grouping, where our model is forced to distinguish the positive pairs from their corresponding negatives, which is demonstrated to be helpful in experiments. We also conduct a detailed analysis of the effectiveness of this method in Section 4.5. The training loss is the same as the original, where cross-entropy loss is used to maximize the diagonal elements and minimize all other entries in the similarity matrix.
For the generative models like LLaVA Liu et al. (2023b, a), we reformat our data into conversations and follow their exact training paradigm to fine-tune, as shown in Figure 3 (b). The training loss is the original next-token prediction loss.
@@ 4 Experiments
We utilize the proposed three datasets, Flickr30k-Positions, Flickr30k-Counting, and Flickr30k-Attributes, created via CounterCurate, to fine-tune both contrastive and generative models. Specifically, we select two common multimodal models, ViT-B/32 from Ilharco et al. (2021) OpenCLIP pre-trained on LAION-2B Schuhmann et al. (2022) and LLaVA-1.5 Liu et al. (2023a) as base models. An overview of our results is attached in Figure 5.
@@ 4.1 Implementation Details
For Flickr30k-Attributes, we trained the CLIP model with a learning rate of 1e-5, batch size of 256, and mixed precision for 5 epochs. For LLaVA-1.5, we used a learning rate of 2e-6 and batch size of 16 for 1 epoch. For Flickr30k-Positions, we used a learning rate of 2.56e-5 and batch size of 1024 for CLIP, and trained for 50 epochs without grouping and 15 epochs with grouping. For Flickr30k-Counting, CLIP was trained with a learning rate of 5e-5. The LLaVA parameters for both Flickr30k-Positions and Flickr30k-Counting were identical, with a learning rate of 2.56e-5 and batch size of 16.

### v3

@@ Abstract
We first spotlight the near-chance performance of multimodal models like CLIP and LLaVA in physically grounded compositional reasoning. We then apply simple data augmentation using the grounded image generation model GLIGEN to generate fine-tuning data, resulting in significant performance improvements: +33% and +37% for CLIP and LLaVA, respectively, on our newly curated Flickr30k-Positions benchmark. Moreover, we exploit the capabilities of high-performing text generation and image generation models, specifically GPT-4V and DALLE-3, to curate challenging semantic counterfactuals, thereby further enhancing compositional reasoning capabilities on benchmarks such as SugarCrepe, where CounterCurate outperforms GPT-4V.
To facilitate future research, we release our code, dataset, benchmark, and checkpoints at https://countercurate.github.io/.
@@
Jianrui Zhangthanks: Equal Contribution.1  Mu Cai∗1  Tengyang Xie1,2  Yong Jae Lee1 {harrisz,mucai,tx,yongjaelee}@cs.wisc.edu 1University of Wisconsin–Madison      2Microsoft Research https://countercurate.github.io
@@ 1 Introduction
Currently, the majority of research focuses on semantic compositional reasoning, leaving another fundamental problem, physically grounded compositional reasoning, under-explored. This involves tasks such as counting and distinguishing left/right and up/down positions between objects. For example, in Figure 1, GPT-4V OpenAI (2023c) made the wrong choice even though it identified all objects in question, demonstrating capable semantical yet weak physical compositional reasoning. [^1] We were able to reproduce our findings of GPT’s weakness in multiple fashions: the failure cases themselves were initially discovered via manual testing before the APIs were available; we were able to both use the 2023-07-01 Azure API and the website version of GPT4-V to recreate many of the failure cases as well.
@@
With such an observation, it is not surprising to find that both contrastive models like CLIP Radford et al. (2021) and generative models like LLaVA Liu et al. (2023b) deliver near random chance performance on our newly curated benchmarks. We hypothesize that modern LMMs are largely oblivious to positional differences. To address this, we generate counterfactual image-text pairs using simple methods, such as horizontal flipping for left/right distinctions, as well as by leveraging a grounded image generation model, GLIGEN Li et al. (2023), to accurately replace/remove objects of interest for up/down and counting tasks. Our approach shows significant improvements such as 33+for CLIP and 37+for LLaVA on our Flickr30k-Positions benchmark.
Existing methods also do not fully utilize the capabilities of powerful generative models when creating semantic counterfactual fine-tuning data. We argue that the key to enhancing compositional reasoning capabilities lies in the careful curation of accurate and sufficiently challenging image and text counterfactuals. For example, augmented negative captions with linguistic rules used by recent fine-tuning approaches (e.g., Yuksekgonul et al., 2023; Le et al., 2023) can unintentionally follow unnatural language distributions, which are easily distinguishable by a text-only model without any image evidence (Lin et al., 2023).
@@ 2.2 Compositional Reasoning
NegCLIP Yuksekgonul et al. (2023) attempts to improve compositional reasoning via fine-tuning CLIP with perturbed captions as the negative text samples in contrastive learning. However, such permuted captions can be easily detected by a text-only transformer Lin et al. (2023), demonstrating that a more effective approach is needed to further enhance compositional reasoning. DAC Doveh et al. (2024) leverages dense and aligned captions for better positive captions. SGVL Herzig et al. (2023) utilizes scene graphs to fine-tune the pretrained vision-language model. TSVLC Doveh et al. (2023) leverages language models to generate positive and negative captions. COCO-Counterfactual Le et al. (2023) further utilizes both negative images and negative captions to fine-tune CLIP via a rule-based data curation pipeline.
Moreover, previous works mainly consider composition reasoning with semantically different concepts rather than physically grounded relationships such as object counting and distinguishing left/right and above/below. Though a recent work Paiss et al. (2023) improved CLIP’s counting capabilities, they neither addressed compositional reasoning nor leveraged negative images. In our paper, we systematically improve both semantic and physical compositional reasoning by leveraging the most capable generative models including GPT-4V OpenAI (2023c) and DALLE-3 Betker et al. (2023) to generate counterfactual image/text pairs.
@@ 2.3 Counterfactual Reasoning
Figure 3: Fine-tuning different types of large multimodal models with CounterCurate. Our pipeline can enhance both contrastive learning models and generative models by augmenting vanilla image-caption pairs with curated negative images and captions. Specifically, our counterfactual image-caption pairs (a) provide auxiliary contrastive loss for models like CLIP, where positive contrastive units in the similarity matrix are colored as blue/red and negative ones are colored as white, and (b) can be naturally integrated into the original next-token prediction loss in text generation models such as LLaVA.
Counterfactual reasoning Morgan and Winship (2015) refers to the process of imagining “what if” scenarios by modifying the input data. In the context of visio-linguistics scenarios, this involves curating negative images and captions by manipulating the original data and observing how it affects the outcome. This helps models understand cause and effect and predict outcomes in situations they’ve never seen before. The key to counterfactual reasoning is curating meaningful and hard enough examples. COCO-Counterfactual Le et al. (2023) explores simple linguistic rules to generate negative captions and uses an image editing model Prompt2Prompt Hertz et al. (2023) to produce negative images Hertz et al. (2023). Prompt2Prompt is less desirable in understanding complex language instructions, as the generated counterfactuals are less reliable and challenging.
In this paper, we collect the very challenging negative image and text examples by leveraging the most capable language models GPT-4 OpenAI (2023b), text-to-image generation model DALLE-3 Betker et al. (2023), and GLIGEN Li et al. (2023), significantly improving our model’s reasoning. Utilizing these hard negatives, our models can better grasp the nuances of language and vision, enhancing their performance on tasks that require a sophisticated understanding of the world.
@@ 3 Approach
In Sec. 3.1, we first explain our selection of the Flickr30k Entities dataset for both benchmarking and fine-tuning. Then in Sec. 3.2 and 3.3, we demonstrate the near-chance performance of multimodal models like CLIP and LLaVA on physically grounded compositional reasoning tasks, including positional understanding and counting, and improve their performance by generating negative image-text pairs via data augmentation and grounded image inpainting. Finally, we introduce using capable text and image generation models to improve models’ semantic compositional reasoning in Sec. 3.4.
@@ 3.1 Necessity of Grounded Image Captions
Previous approaches Yuksekgonul et al. (2023); Hsieh et al. (2023) adopt global image captioned datasets such as COCO-Captions Chen et al. (2015) and craft negative captions to either benchmark or improve compositional reasoning. This method, however, lacks regional information important for physically grounded reasoning, such as determining the relative position of two objects. Therefore, we leverage Flickr30k Entities Plummer et al. (2016), a grounded image captioning dataset to both curate negative image-caption pairs and benchmark models.
@@ 3.2 Improving Positional Understanding
in Flickr30k Entities Plummer et al. (2016) for each object $j$ in Image $I$, where box $B_{j}$’s upper-left and lower-right corners are at $(x_{j1},y_{j1})$ and $(x_{j2},y_{j2})$, respectively. In order to ensure non-ambiguity in relative position descriptions between categories (e.g., “the dog is to the left of the cat”), we only consider images where for each object category, there is only at most one instance of it in the image. For each pair of objects $\{a,b\}\subseteq I$, we identify the positional object pairs via the formula:
$\begin{cases}x_{a2}\leq x_{b1}\implies\text{ $a$ is to the {left} of $b$}\\ x_{a1}\geq x_{b2}\implies\text{ $a$ is to the {right} of $b$}\\ y_{a2}\leq y_{b1}\implies\text{ $a$ is {above} $b$}\\ y_{a1}\geq y_{b2}\implies\text{ $a$ is {below} $b$}\\ \end{cases}$
@@
We also prompt text-only GPT4 OpenAI (2023b) to rewrite the vanilla negative prompt $a$ is above/below $b$ to better guide image generation models. For example, the caption “a street is above a young child" is unnatural, while the GPT-expanded caption “the street stretches out above a young child, elevated by a bridge” is much better. For detailed prompt engineering, see Appendix B.
@@
Furthermore, we include negative images in our dataset to curate Flickr30k-Positions. For the left-right pairs, we apply a simple data augmentation method: flip the original images horizontally. For the above-below pairs, however, flipping vertically in most cases would lead to unreasonable images. To address this issue, we utilize GLIGEN Li et al. (2023), an object-level text-to-image generation model that can ground text prompts with bounding boxes. In particular, for each above-below object pair $(a,b)\in I$, we swap the centers of the bounding boxes $B_{a},B_{b}$ while maintaining their widths and heights, ensuring that the generated images are natural while preserving the objects’ sizes and aspect-ratios as demonstrated in Figure 4. The new bounding boxes and prompts are fed to GLIGEN to generate the corresponding negative image $I^{\prime}$.
@@ 3.3 Improving Object Counting
If all instances of the two selected object categories do not spatially overlap with any other object in the image, we generate the corresponding negative caption by decrementing the one with more objects, $Q$, and incrementing the one with fewer, $P$; for example, $C^{\prime}=$“there are four cats and three dogs”. This enforces a hard counterfactual format as the numbers in the positives are reversed in the negatives.
In most cases, however, objects in an image do overlap, and we cannot ascertain that the increment/decrement rule will still generate hard counterfactuals. For example, if a dog overlaps with a cat, and we re-use the rule above to remove one dog and add one cat, we end up with “three cats and three dogs”. In this case, we only remove either one of $P$ or one of $Q$, for example, the dog, and all objects it overlaps with. An example of the resulting negative caption is $C^{\prime}=$“there are two cats and three dogs.”
@@
In the case where objects do not overlap, we simply leverage GLIGEN Li et al. (2023) to inpaint a new object of the incremented category $P$ at the bounding box of one of the objects from the decremented category $Q$. Otherwise, we remove the target objects using GLIGEN to replace it with an object from a generic category. We simply use the plant category for replacement, as plants can appear in almost any natural setting, let it be outdoor or indoor, and we find that GLIGEN does well in generating it according to each image $I$’s background during inpainting for each corresponding negative image $I^{\prime}$. More details can be found in Appendix C.
@@ 3.4 Improving Semantic Compositional Reasoning
Existing works Yuksekgonul et al. (2023); Le et al. (2023) use a rule-based approach to generate negative images/captions for semantic compositional reasoning, resulting in less desirable negatives such as uncommon negative captions e.g., “A dog with blue fur" from “A dog with black fur". We introduce Flickr30k-Attributes as in Figure 2 (c), which utilizes GPT-4V OpenAI (2023c) to generate hard negative captions and leverages DALLE-3 Betker et al. (2023) for generating corresponding high-quality negative images. We extract the first and most detailed caption $C$ for each image $I$ from the original Flickr30k dataset Young et al. (2014).
@@
To generate high-quality negative captions, we leverage GPT-4V and feed both positive images $I$ and captions $C$ to generate negative captions $C^{\prime}$. To let GPT-4V know the objects’ locations, we overlay bounding boxes $B$ onto the original positive image Cai et al. (2024) and also feed this annotated image into GPT-4V. We prompt GPT-4V to generate three kinds of captions, including (i) changing a noun, (ii) changing an adjective, and (iii) swapping any of the two phrases, nouns, or adjectives. For example, given the original caption $C$ “a man wearing a black shirt and blue jeans", GPT-4V can generate the following negatives $C^{\prime}$: (i) “a man wearing a black jacket and blue jeans", (ii) “a man wearing a black shirt and red jeans", and (iii) “a man wearing a blue shirt and black jeans". Detailed prompts to GPT-4V are shown in Appendix A.
@@ 3.5 Fine-tuning Methodology
As demonstrated in Figure 3, CounterCurate can be used to fine-tune both contrastive Radford et al. (2021) and generative Liu et al. (2023b) multimodal models. To fine-tune a contrastive model like CLIP, we sample $N$ positive image-text pairs and the corresponding negative image-text pairs to form a batch $(C,I,C^{\prime},I^{\prime})$ as shown in Figure 3 (a). We refer to this method as Grouping, where the model is forced to distinguish the positive pairs from their corresponding negatives, which is demonstrated to be helpful in our experiments. We also conduct a detailed analysis of the effectiveness of this method in Section 4.5. The training loss is the same as the original, where cross-entropy loss is used to maximize the diagonal elements and minimize all other entries in the similarity matrix.
For generative models like LLaVA Liu et al. (2023b, a, 2024), we reformat our data into conversations and follow their exact training paradigm to fine-tune, as shown in Figure 3 (b). The training loss is the original next-token prediction loss.
@@ 4 Experiments
We utilize the proposed three datasets, Flickr30k-Positions, Flickr30k-Counting, and Flickr30k-Attributes, created via CounterCurate, to fine-tune both contrastive and generative models. Specifically, we select two common multimodal models, ViT-B/32 from Ilharco et al. (2021) OpenCLIP pre-trained on LAION-2B Schuhmann et al. (2022) and LLaVA-1.5 Liu et al. (2023a) as base models. An overview of our results is shown in Figure 5.
@@ 4.1 Implementation Details
For Flickr30k-Attributes, we train the CLIP model with a learning rate of 1e-5, batch size of 256, and mixed precision for 5 epochs. For LLaVA-1.5, we use a learning rate of 2e-6 and batch size of 16 for 1 epoch. For Flickr30k-Positions, we use a learning rate of 2.56e-5 and batch size of 1024 for CLIP, and train for 50 epochs without grouping and 15 epochs with grouping. For Flickr30k-Counting, we train CLIP with a learning rate of 5e-5. The LLaVA parameters for both Flickr30k-Positions and Flickr30k-Counting are identical, with a learning rate of 2.56e-5 and batch size of 16.